OpenStudy (just_gymnast_me):

Can anyone tell me why x^2 − 8x − 11 is prime? I know it is prime, I just don't know how to explain why it is.

OpenStudy (just_gymnast_me):

\[x ^2 - 8 - 11\]

OpenStudy (just_gymnast_me):

is the trinomial

OpenStudy (bibby):

a number is prime if it can only be divide by 1 and itself, right?

OpenStudy (just_gymnast_me):

yes

OpenStudy (just_gymnast_me):

so if one part in the equation is prime the equation itself is prime?

OpenStudy (bibby):

I guess this equation has no whole number factors

OpenStudy (bibby):

or solutions*

OpenStudy (just_gymnast_me):

im confused now :?

OpenStudy (bibby):

ok what are the solutions of \(x^2-16=0\)

OpenStudy (just_gymnast_me):

(x+4) (x-4)

OpenStudy (bibby):

you can't factor it like that because you can't write -8 in terms of the factors of 11, you'd have to complete the square

OpenStudy (bibby):

\(x^2 − 8x − 11=(x-4)^2-27\)

OpenStudy (just_gymnast_me):

Then how do i make x2 − 8x − 11 a prime equation?

OpenStudy (just_gymnast_me):

its asking to find a prime trinomial and explain why it is prime. I went to another site and the said it was prime so i came here. So i need a new trinomial...

OpenStudy (bibby):

it is prime

OpenStudy (bibby):

http://www.saylor.org/site/wp-content/uploads/2013/02/RWM102-9.3.4-Prime-Trinomials-FINAL.pdf remember what I said about the factors of 11?

OpenStudy (just_gymnast_me):

yes

OpenStudy (just_gymnast_me):

only 11 and 1 because it is prime

OpenStudy (bibby):

\(x^2-8x-11\) we want factors of -11 that add/subtract up to -8 Listing the factors of -11: -1 x 11= -11 none (well in this case we only have one) of these add up to -8 so this can't be factored

OpenStudy (just_gymnast_me):

that makes sense now

OpenStudy (just_gymnast_me):

thank you (:

OpenStudy (bibby):

anytime c: